Cherry-red embers in a fireplace are at 835°C and have an exposed area of 0.240 m2 and an emissivity of 0.980. The surrounding room has a temperature of 15.0°C. If 50% of the radiant energy enters the room, what is the net rate of radiant heat transfer in kilowatts?

Radiation makes it impossible to stand close to a hot lava flow. Calculate the rate of heat transfer by radiation in kilowatts from 1.00 m2 of 1140°C fresh lava into 35.0°C surroundings, assuming lava's emissivity is 1.
Answer #1

rate of heat tranfer H = Q/t = A e Sigma T^4

here A is area

e is emissivity

Sgma is Stefans constant

T is absolute temperature

H   = Q/t = 0.24 * 0.98 * 5.67 *10^-8 *(835 + 273)^4

rate of heat transfer = 20 kWatts

50% of this = 10 k Watts

--------------------------------------

again

H = 1 * 1 * 5.67 *10^-8 * ((1140+273)^4 - (35+273)^4)

H = 225.5 kWatts